5. Suppose a monopolist has TC = 100 + 10Q + 2Q2, and the demand curve it faces is P = 90 - 2Q. What will be the price, quantity, and profit for this firm?

1
Expert's answer
2022-01-10T09:55:50-0500

"TC = 100 + 10 Q + 2Q^2 \\

MC = 10 + 4Q \\

P = 90 - 2Q \\

TR = PQ \\

= 90Q - 2Q^2 \\

MR = 90 - 4Q"

At equilibrium MR = MC

"90 - 4Q = 10 + 4Q \\

8Q = 80 \\

Q = \frac{80 }{ 8} = 10"

(equilibrium quantity)

"Price = 90 - 2Q \\

= 90 - 2 \times 10 \\

= 90 - 20 \\

= 70"

(equilibrium price)

"Profit = TR - TC \\

= P \times Q - (100 + 10 \times 10 + 2\times 10^2 ) \\

= (70 \times 10 ) - (100 + 10 \times 10 + 2 \times 10^2 ) \\

= 700 - ( 100 + 100 + 2 \times 100 ) \\

= 700 - ( 100 + 100 + 200 ) \\

= 700 - 400 = 300"

Answer: Price = 70 , Quantity = 10 , Profit = 300